Chapter 2 Assignment X M MHE Reader x + - X € → C newconnect.mheducation.com/flow/connect.html Chapter 2 Assignment Saved Hel2 Chapter 2 Assignment X M MHE Reader x + € → C newconnect.mheducation.com/flow/connect.html Chapter 2 Assignment Help Save &2 Chapter 2 Assignment X M MHE Reader x + € → C newconnect.mheducation.com/flow/connect.html Chapter 2 Assignment Help Save &I only need help with part i/10
Answer #1

Solution:

Actual Overhead incurred = 20200+62400+22800+19600+5900 = $130,900

Applied overhead = 14383*$11 = $158,213

Overapplied Overhead = 158,213 - 130,900 = $27,313

Entry 10 (i):

Manufacturing Overhead Debit............$27313

Cost of Goods sold...............................................$27,313